A rock starts at ground level with 741 joules of kinetic energy and rises until it ceases moving. What is its gravitational potential energy at this point?
1
Expert's answer
2016-03-18T15:48:05-0400

Since the rock starts at ground level, its initial gravitational potential energy is zero.
At the same time, its total energy Etotal = EP + EK = 0 + 741 = 741 J
When the rock moves upwards, its kinetic energy is transformed into gravitational energy. At the highest point of the trajectory, the rock ceases moving, therefore its kinetic energy is zero. At the same time, the rock's total energy remains constant due to the energy conservation: Etotal = PE + PK = const.
At the highest point, PE = Etotal - EK = 741 - 0 = 741 J.
It means that the rock's initial kinetic energy has completely transformed into the gravitational potential energy.